Jenny's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Jenny

$4.05
per pound, and type B coffee costs

$5.45
per pound. This month's blend used three times as many pounds of type B coffee as type A, for a total cost of

$652.80
. How many pounds of type A coffee were used?

Answer by stanbon(75887)

Jenny's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Jenny $4.05 per pound, and type B coffee costs $5.45 per pound. This month's blend used three times as many pounds of type B coffee as type A, for a total cost of $652.80
. How many pounds of type A coffee were used?
------------
Equations:
B = 3A
4.05A + 5.45B = 652.80
------------------------
Substitute for "B" and solve for "A"::
405A + 545*3A = 65280
---
405A + 1635A = 65280
2040A = 65280
A = 32 lbs (amt. of A coffee used)
------------
Cheers,
Stan H.
-------------